OpenStudy (anonymous):

What is the experimental probability that exactly 3children in a family of 4 children will be boys? Assume that P(boy)=P(girl)

OpenStudy (raden):

the number events of (B,B,B,G) = 4!/3! = 4 ways P(B) = 1/2 and P(G) = 1/2 so, P(B,B,B,G) = (1/2 * 1/2 * 1/2 * 1/2) * 4 = 1/4 assumed this is an independency event